OpenStudy (anonymous):

Give exact and approximate solutions to three decimal places. (x-6)^2=64 Thanks

OpenStudy (anonymous):

(x-6)^2 = 64 taking squareroot of both sides (x-6) = +/- 8 x = +/-8 + 6 so x = 8+ 6 = 14 or x = -8 + 6 = -2